# Infer-and-widen versus split-and-condition: two tales of selective inference

### v1

@@ 5 Vignette #3: Inference after the lasso
We now consider inference on the set of coefficients selected by the lasso in a linear model (Tibshirani, 1996). We again consider the model 3.1, alongside a fixed design matrix $X\in\mathbb{R}^{n\times p}$ with unit-normalized columns, i.e. $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Our new selection rule is the magnitude-ordered set of features selected by the lasso:
$\bar{\mathcal{S}}(Y)=\left\{j_{1},\dots,j_{l}:\left|\hat{\beta}_{j_{1}}^{\lambda}\right|>\dots>\left|\hat{\beta}_{j_{l}}^{\lambda}\right|>0,\;\underset{k\notin\{j_{1},\ldots,j_{l}\}}{\max}|\hat{\beta}_{k}^{\lambda}|=0\right\}$ (5.1)
$\mathrm{where}\quad\hat{\beta}^{\lambda}:=\mathop{\mathrm{arg\,min}}_{\beta}\left\{\left\|Y-X\beta\right\|^{2}_{2}+\lambda\|\beta\|_{1}\right\}.$
We then consider a least squares model fit to the features in $\bar{\mathcal{S}}(Y)$. The target of inference is the coefficient of the feature with largest lasso magnitude, i.e. $\beta_{{\bar{\mathcal{S}}(Y)},1}$ where
$\beta_{\bar{\mathcal{S}}(Y)}:=\left(X_{\bar{\mathcal{S}}(Y)}^{\top}X_{\bar{\mathcal{S}}(Y)}\right)^{-1}X_{\bar{\mathcal{S}}(Y)}^{\top}\mathbb{E}\left[Y\right].$ (5.2)
Here, the design matrix $X_{\bar{\mathcal{S}}(Y)}$ has columns $X_{j_{1}},\dots,X_{j_{l}}$ ordered by the selected indices $\bar{\mathcal{S}}(Y)$.
Zrnic and Jordan (2023) do not propose an infer-and-widen interval based on the lasso algorithm, as the lasso is not stable in the sense of Definition 1: while they do consider a variant of the lasso that is stable, it is far enough from the original lasso proposal that we do not investigate it here. Instead, we construct a split-and-condition interval based on Gaussian data thinning (Neufeld et al., 2024; Dharamshi et al., 2024), and then will consider the “oracle” infer-and-widen interval that makes use of the identical selection event.
Towards this end, we define
@@
Under the model (3.1), consider the selection rule $\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(\cdot)$ defined in (5.3). A $1-\alpha$ split-and-condition interval for $\beta_{{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)},1}$ , where
$\beta_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}:=\left(X_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}^{\top}X_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}\right)^{-1}X_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}^{\top}\mathbb{E}\left[Y\right],$
takes the form
@@
where
$\hat{\beta}_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}:=\left(X_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}^{\top}X_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}\right)^{-1}X_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}^{\top}\left(Y-\tfrac{1}{c}\zeta\right),$
@@
We compare the widths of the split-and-condition (5.4) and “oracle” infer-and-widen intervals — both of which have identical selection events — in Figure 3. We simulate data according to 3.1 with $n=100$ and $\sigma=1$. We set $\mu=X^{\top}\phi$, where half of the elements of $\phi$ are zero and the other half are either zero or drawn from exponential distributions with means $50/7$ or $5$ in panels (a), (b), and (c) of Figure 3. The rows of $X$ are independent draws from ${\bf{\rm N}}\,_{p}(0,\rho(1_{p}1_{p}^{\top}-I_{p})+I_{p})$ with $\rho=0.5$, and the columns of $X$ are scaled so that $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Figure 3 displays the ratio of confidence interval widths as we vary $p$ and the variance of the Gaussian noise $c^{2}$ in (5.3).
We see from Figure 3 that the relative widths of the two intervals depend on details of the noise variance in (5.3) and distribution of $\phi$. Of course, the “oracle” infer-and-widen interval cannot be computed in practice: any real-world infer-and-widen interval that achieves nominal coverage would necessarily be at least as wide as the oracle interval, if not far wider. Thus, while it may be possible to develop an infer-and-widen interval that might outperform the simple split-and-condition interval in 5.4 in specific scenarios, the potential upside of any such approach is extremely limited.
Figure 3: Vignette #3, under the model (3.1) with $n=100$, $\sigma=1$, $\mu=X^{\top}\phi$, and selection rule in 5.3. The columns of $X$ are independent normal draws with column correlation $0.5$. Results are averaged over $1000$ simulated datasets. We consider inference on a lasso-selected coefficient in a least squares model, varying the mean of $\phi$ across $0,5,50/7$ in panels (a), (b), and (c), respectively. We display the ratio of widths of an “oracle” infer-and-widen interval to the split-and-condition interval (5.4), with identical selection events; a ratio above one indicates that the split-and-condition interval is narrower and hence preferable. In many settings, even the unattainable “oracle” infer-and-widen interval is wider than the easy-to-compute split-and-condition interval.
@@ 6 The bias of infer-and-widen intervals
We have seen in the previous sections that when infer-and-widen and split-and-condition intervals are formed for the same data-driven target of inference, the former tends to yield intervals that are far wider than the latter. We now investigate the reason for this. To facilitate the analysis, we consider Vignette #1 with a simple selection rule that uses Gaussian noise (as opposed to the Laplace noise used in (3.3)),
@@
We know from Section 2.2 that any infer-and-widen interval is centered around $Y_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}$.
@@
Under (3.1) with $\mu=0$, consider the selection rule $\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)$ in (6.1). Then
$\mathbb{E}\left[Y_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}-\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\right]\geq\frac{\sigma}{2}(\sqrt{1+c^{2}}-c)\sqrt{\log n}.$ (6.2)
That is, the bias increases on the order of $\sqrt{\log n}$. Note that as $c\to\infty$, i.e. the selection rule becomes fully random, this lower bound goes to zero as expected.
Noting that $\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)$ in (6.1) is independent of $Y-\frac{1}{c}\zeta$, it is straightforward to show that $[(Y-\frac{1}{c}\zeta)_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\pm\sigma(1+c^{-2})^{1/2}z_{1-\alpha/2}]$ is a valid $1-\alpha$ split-and-condition confidence interval for $\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}$. Furthermore, $\mathbb{E}\left[\left(Y-\frac{1}{c}\zeta\right)_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}-\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\right]=0$, i.e. the midpoint of the split-and-condition interval is unbiased for the mean of the selected observation.
Thus, for finite $n$, infer-and-widen here leads to a confidence interval that is centered in the wrong place, and the situation worsens as $n$ increases. By contrast, split-and-condition leads to a confidence interval that is correctly centered.
@@ 7 Discussion
Towards a better systematic understanding of the plethora of methods available for post-selection inference, in this paper we classified these methods into one of two frameworks: infer-and-widen and split-and-condition. We investigated the performance of these frameworks in three vignettes: two taken from Zrnic and Jordan (2023), and the third involving feature selection via the lasso. We constructed split-and-condition intervals that had exactly the same selection events as recently-proposed infer-and-widen intervals (Vignettes #1 and #2) or as an “oracle” infer-and-widen approach. Our findings are as follows:
@@
    In the first two vignettes, when we devise split-and-condition and infer-and-widen intervals that have identical selection events, the former are substantially narrower than the latter.
@@
    For a given nominal coverage, the “oracle” infer-and-widen interval is substantially narrower than the interval arising from algorithmic stability, which is the current state-of-the-art approach for the infer-and-widen framework (Zrnic and Jordan, 2023). Stated another way, while algorithmic stability may be the best infer-and-widen method to date, there is still room for improvement in constructing tighter intervals.
@@
    Nonetheless, we can in some cases obtain a simple and easy-to-compute split-and-condition interval that outperforms the “oracle” infer-and-widen approach (in the sense of having a narrower interval width for a given selection error). Thus, the benefit of developing improved infer-and-widen methods is limited or non-existent.
Why does split-and-condition so often outperform infer-and-widen? The answer is simple: the latter yields intervals centered around a biased estimator of the target of inference, whereas the former faces no such restriction. This point was illustrated in Section 6.
We close with two remarks, firstly about the case of a deterministic selection rule. Zrnic and Jordan (2023) point out that because such a rule is $(0,0,0)$-stable in the sense of Definition 1, the algorithmic stability interval is simply the classical interval (see Theorem 2). Of course, the same is true for split-and-condition: in the case of a deterministic selection rule, we can take $Y^{\mathrm{train}}=\varnothing$ and $C(Y)=S(Y^{\mathrm{train}})=\varnothing$, using the notation in Section 2; then, the split-and-condition interval is simply the classical interval.
Secondly, it is worth mentioning the results of Goeman and Solari (2022), who proved that any split-and-condition method is dominated by an unconditional multiple testing procedure defined on the full universe of hypotheses. In contrast, we examine unconditional methods which (1) are defined only on the selected hypotheses and (2) are in the infer-and-widen framework. For instance, a Bonferroni correction provides valid inference on the selected mean in Vignette #1, but is dominated in power by Bonferroni-corrected tests of all possible means. Our comparisons reflects a study of existing approaches and practices.
@@ Acknowledgements
Appendix A Vignette #1
A.1 Proof of Proposition 2
To construct a split-and-condition confidence interval under the selection rule 3.2, it suffices to characterize the distribution $Y\mid Y^{\mathrm{train}}$, where $Y^{\mathrm{train}}:=Y+\zeta$ and $\zeta_{1},\ldots,\zeta_{n}\stackrel{{\scriptstyle\mathrm{ind.}}}{{\sim}}\mathrm{Laplace}\left(c\right)$. This conditional distribution is difficult to characterize, however, and so we instead consider the distribution conditional on $Y^{\mathrm{train}}$ and an additional event.
@@ A.1 Proof of Proposition 2
$Y\mid(Y^{\mathrm{train}},\delta)\sim{\bf{\rm N}}\,_{n,A}(\mu+c^{-1}\sigma^{2}\delta,\sigma^{2}I_{n}),$ (A.1)
@@
$\propto\mathbb{I}[Y\in A]\exp\left\{-\frac{1}{2\sigma^{2}}(Y-\mu-c^{-1}\sigma^{2}\delta)^{\top}(Y-\mu-c^{-1}\sigma^{2}\delta)\right\}.$ (A.2)
@@
A.2 Proof of Proposition 1
First, since $\mu=0$,

### v2

@@ 5 Vignette #3: Inference after the lasso
We now consider inference on the set of coefficients selected by the lasso in a linear model (Tibshirani, 1996). We again consider the model 3.1, alongside a fixed design matrix $X\in\mathbb{R}^{n\times p}$ with unit-normalized columns, i.e., $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Our new selection rule is the magnitude-ordered set of features selected by the lasso:
$\bar{\mathcal{S}}(Y)=$$\bigg{\{}j_{1},\dots,j_{l}:\left|\hat{\beta}_{j_{1}}^{\lambda}\right|>\dots>\left|\hat{\beta}_{j_{l}}^{\lambda}\right|>0,\;$ (5.1)
$\quad\underset{k\notin\{j_{1},\ldots,j_{l}\}}{\max}|\hat{\beta}_{k}^{\lambda}|=0\bigg{\}},$
where
$\hat{\beta}^{\lambda}:=\mathop{\mathrm{arg\,min}}_{\beta}\left\{\left\|Y-X\beta\right\|^{2}_{2}+\lambda\|\beta\|_{1}\right\}.$
We then consider a least squares model fit to the features in $\bar{\mathcal{S}}(Y)$. Our target of inference is the coefficient of the feature with largest lasso magnitude, i.e., $\beta_{{\bar{\mathcal{S}}(Y)},1}$, where
$\beta_{\bar{\mathcal{S}}(Y)}:=A_{\bar{\mathcal{S}}(Y)}\mathbb{E}\left[Y\right]$
and
$A_{J}:=\left(X_{J}^{\top}X_{J}\right)^{-1}X_{J}^{\top}$
for any subset of features $J\subseteq[p]$.
The lasso is not stable in the sense of (Zrnic and Jordan, 2023). Therefore, Zrnic and Jordan (2023) consider a variant of the lasso that is far enough from the original lasso proposal that we do not investigate it here. Instead, in what follows, we construct an interval based on Gaussian data thinning (Neufeld et al., 2024; Dharamshi et al., 2024) that yields conditional guarantees in the sense of 1.3, and then consider the “oracle” infer-and-widen interval that makes use of the identical selection event.
Towards this end, for some $c>0$, we define
@@
Under (3.1), consider the selection rule $\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(\cdot)$ defined in (5.2). The data thinning confidence interval
@@
has $1-\alpha$ conditional selective coverage for $\beta_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y),1}$ in the sense of 1.3 with $\mathcal{C}(Y):=\left(Y+c\zeta\right),$ where
$\hat{\beta}_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}:=A_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}\left(Y-\tfrac{1}{c}\zeta\right)$
and
@@
We compare the widths of the data thinning (DT) (5.3) and “oracle” infer-and-widen intervals — with identical selection events — in Figure 3. We simulate data according to 3.1 with $n=100$ and $\sigma=1$. We set $\mu=X\phi$, where half of the elements of $\phi$ are zero and the other half are either zero or drawn from exponential distributions with means $50/7$ or $5$ in panels (a), (b), and (c) of Figure 3. The rows of $X$ are independent draws from ${\bf{\rm N}}\,_{p}(0,\rho(1_{p}1_{p}^{\top}-I_{p})+I_{p})$ with $\rho=0.5$, and the columns of $X$ are scaled so that $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. The penalty $\lambda$ is adaptively chosen using 3-fold cross-validation. Figure 3 displays the ratio of the average confidence interval widths as we vary $p$ and the variance of the Gaussian noise $c^{2}$ in (5.2).
We see from Figure 3 that the relative widths of the two intervals depend on the variance of the noise in (5.2) and distribution of $\phi$. Of course, the “oracle” infer-and-widen interval cannot be computed in practice: any real-world infer-and-widen interval that achieves nominal coverage would necessarily be at least as wide as the oracle interval, if not far wider. Thus, while it may be possible to develop an infer-and-widen interval that might outperform the simple data thinning interval in 5.3 in specific scenarios, the potential upside of any such approach is extremely limited.
Figure 3: Vignette #3 under (3.1) with $n=100$, $\sigma=1$, $\mu=X\phi$, and the selection rule in 5.2. The columns of $X$ are independent normal draws with column correlation $0.5$. Results are averaged over $1000$ simulated datasets. We consider inference on a lasso-selected coefficient in a least squares model, varying the mean of $\phi$ across $0,5,50/7$ in panels (a), (b), and (c), respectively. We display the ratio of average widths of the 95% “oracle” infer-and-widen interval and the 95% data thinning (DT) interval (5.3), with identical selection events and level $\alpha=0.05$; a positive log ratio indicates that the data thinning interval is narrower and hence preferable. In many settings, even the unattainable “oracle” infer-and-widen interval is wider than the attainable data thinning interval.
@@ 6 The bias of infer-and-widen intervals
We now investigate why infer-and-widen intervals are too wide. To facilitate the analysis, we consider Vignette #1 with a simple selection rule that uses Gaussian noise,
@@
We know from Section 2 that any infer-and-widen interval is centered around $Y_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}$.
@@
Under (3.1), with $\mu_{1}=\dots=\mu_{n}$, consider the selection rule $\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)$ in (6.1). Then
$\mathbb{E}\left[Y_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}-\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\right]\geq\frac{\sigma}{2}(\sqrt{1+c^{2}}-c)\sqrt{\log n}.$
That is, the bias increases on the order of $\sqrt{\log n}$. Note that as $c\to\infty$, i.e., as the selection rule becomes fully random, this lower bound goes to zero as expected.
Noting that $\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)$ in (6.1) is independent of $Y-\frac{1}{c}\zeta$ (Neufeld et al., 2024), it is straightforward to show that $[(Y-\frac{1}{c}\zeta)_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\pm\sigma(1+c^{-2})^{1/2}z_{1-\alpha/2}]$ is a valid $1-\alpha$ selective confidence interval for $\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}$. Furthermore,
$\mathbb{E}\left[\left(Y-\frac{1}{c}\zeta\right)_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}-\mu_{\mathcal{S}_{c}^{\mathrm{Gaussian}}(Y)}\right]=0,$
i.e., the midpoint is unbiased for the mean of the selected observation.
Thus, infer-and-widen here leads to a confidence interval that is centered in the wrong place, but a simple alternative that makes use of the same selection event yields an interval that is correctly centered.
@@ 7 Discussion
In this paper, we compared the performance of the “infer-and-widen” approach to simple alternative methods in three vignettes: two taken from Zrnic and Jordan (2023), and the third involving feature selection via the lasso. Our findings are as follows:
@@
    In the first two vignettes, infer-and-widen intervals can be computed in practice but are substantially wider than intervals constructed using alternative approaches that make use of identical randomized selection events and yield conditional guarantees.
@@
    In general, for a given nominal coverage, the “oracle” infer-and-widen interval is substantially narrower than an infer-and-widen interval that can be computed in practice.
@@
    In Vignettes #1 and #3, the “oracle” infer-and-widen interval was often wider than the alternative approaches we consider; thus, there may be limited value in developing narrower infer-and-widen methods that can be computed in practice.
One may wonder whether our findings conflict with those of Goeman and Solari (2022), who proved that any method that guarantees conditional coverage (1.3) is dominated by an unconditional multiple testing procedure defined on the full universe of hypotheses. However, their setting is quite different from ours: they consider inference on non-selected parameters and unconditional procedures outside of the infer-and-widen framework.
@@ Acknowledgements
Appendix A A summary of algorithmic stability
The notion of stability underlying the algorithmic stability proposal of [Zrnic and Jordan, 2023] is as follows.
Definition 1 (Definitions 4.1 and 4.2, Zrnic and Jordan [2023]).
Two random variables $Q$ and $T$ are $(\eta,\tau)$-indistinguishable from each other, denoted by $Q\approx_{(\eta,\tau)}T$, if for any measurable set $\mathcal{O}$, it holds that
$\mathbb{P}(Q\in\mathcal{O})\leq e^{\eta}\mathbb{P}(T\in\mathcal{O})+\tau.$ (A.1)
A (possibly randomized) selection rule ${\mathcal{S}}:\mathcal{Y}\mapsto\mathcal{I}$ is $(\eta,\tau,\nu)$-stable with respect to a distribution $\mathcal{P}$ on $\mathcal{Y}$ if there exists a (possibly random, possibly dependent on $\mathcal{P}$) variable ${S}_{0}$ such that
$\mathbb{P}\left(\mathcal{S}(Y)\approx_{(\eta,\tau)}{S}_{0}\right)\geq 1-\nu.$ (A.2)
Since $S_{0}$ is independent of $Y$, we can use the classical interval (1.1) to conduct inference on a target given by $S_{0}$, i.e., it holds that $\mathbb{P}\left(\theta_{S_{0}}\not\in\mathrm{CI}_{S_{0}}^{\alpha}(Y)\right)\leq\alpha$. The key insight underlying algorithmic stability is that if $\mathcal{S}(Y)$ is stable, then for most $Y\sim\mathcal{P}$ the value of $\mathcal{S}(Y)$ is close (in the sense of Definition 1) to $S_{0}$. Thus, it is “almost as if” the selection rule $\mathcal{S}(Y)$ did not involve $Y$. This notion is formalized in Theorem 2.
Theorem 2 (Theorem 5.2, Zrnic and Jordan [2023]).
Let $\mathcal{S}(\cdot)$ be a $(\eta,\tau,\nu)$-stable selection rule with parameters possibly dependent on the data distribution $\mathcal{P}$ of $Y$. Assume that for any fixed $\theta_{i}\in\Theta$, we have access to a confidence interval $\mathrm{CI}_{i}^{\alpha}(Y)$ for $\theta_{i}$ with coverage at least $1-\alpha$, i.e., it satisfies (1.1). Then
$\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{(\alpha-\tau-\nu)e^{-\eta}}(Y)\right)\leq\alpha.$ (A.3)
Thus, to conduct inference on a data-driven target $\theta_{\mathcal{S}(Y)}$, we can apply the infer-and-widen interval in (2.1) with $H(\alpha,{\mathcal{S}})=(\alpha-\tau-\nu)e^{-\eta}$, a function of the stability of the selection rule $\mathcal{S}(\cdot)$. The values of $\eta,\tau,$ and $\nu$ that govern the stability of $\mathcal{S}(\cdot)$ directly determine the width of the confidence interval: the smaller these parameters, the narrower the confidence interval. Zrnic and Jordan [2023] construct stable selection rules in a number of settings via the introduction of auxiliary randomness.
Appendix B Vignette #1
B.1 Proof of Proposition 2
To construct a confidence interval with coverage as in (1.3) under the selection rule (3.2), it suffices to characterize the distribution $Y\mid Y^{\mathrm{train}}$, where $Y^{\mathrm{train}}:=Y+\zeta$ and $\zeta_{1},\ldots,\zeta_{n}\stackrel{{\scriptstyle\mathrm{ind.}}}{{\sim}}\mathrm{Laplace}\left(c\right)$. In the language of Leiner et al. [2023], this is the “data fission” approach. This conditional distribution is difficult to characterize, however, and so we instead consider the distribution conditional on $Y^{\mathrm{train}}$ and an additional event.
@@ B.1 Proof of Proposition 2
$Y\mid(Y^{\mathrm{train}},\delta)\sim{\bf{\rm N}}\,_{n,A}(\mu+c^{-1}\sigma^{2}\delta,\sigma^{2}I_{n}),$ (B.1)
@@
$\propto\mathbb{I}[Y\in A]\exp\left\{-\frac{1}{2\sigma^{2}}(Y-\mu-c^{-1}\sigma^{2}\delta)^{\top}(Y-\mu-c^{-1}\sigma^{2}\delta)\right\}.$ (B.2)
@@
B.2 Proof of Proposition 6
First, without loss of generality, let $\mu_{1}=\dots=\mu_{n}=0$. Then